You want to use someone's pranking ability to predict their ability to sell paper at Dunder Mifflin. You want to use this information to decide on whether or not to higher someone in the future. You have collected pranking ability scores and the number of reams of paper per month on the past 20 sales people that have worked there and found the following regression equation.

b0 = 200

b1 = 20

Questions:

1. Using numbers only, how many reams of paper would you expect someone with zero pranking ability to sell per month?  

2. Let's say Jim had the highest praning ability with a score of 10. How many reams per month (number only) would you expect him to sell?  
Answer #1

The regression equation is,

Number of reams= b0+b1(pranking ability) = 200+20(pranking ability)

1. Number of reams=200+20*0 = 200

2. Number of reams=200+20*10 = 400